0.05 kg of carbon dioxide (molecular weight 44), occupying a volume of 0.03m3 at


1.025bar , is compressed reversibly until the pressure is 6.15bar. calculate the final


temperature the work done on the CO2 and the heat flow to or from the cylinder walls ,


1- when the process is according to the law p.ν1.4 = constant , 2- when the process is


isothermal , assume carbon dioxide to be a perfect gas , and take γ = 1.3

1
Expert's answer
2022-04-24T17:19:10-0400

1.

"A=\frac{p_1V}{\gamma-1}(1-(\frac{p_1}{p_2})^{1-\frac 1{\gamma}})=3.1~kJ,"

2.

"A=p_1V_1\ln\frac {p_1 }{p_2}=5.5~kJ."